if a plane can travel 500 mkiles per hour with the wind and 400 miles per hour against the wind, find the speed of the plane in still air

Answer by jorel1380(3719)

500+400/2=450 mph.
500-450=450-400=50 mph
Still air speed:450 mph. Wind speed:50 mph..